OpenStudy (anonymous):

prove that 3:8 is equivalent to 12:32

OpenStudy (anonymous):

\[3*x=12\] \[8*x=32\] So this means that.. \[12/3=4\] and... \[32/8=4\] So x is equal to 4 Tell me if this helps! (You might need a more in depth explanation, if you do I hope this works)

OpenStudy (jdoe0001):

\(\large { \begin{cases} 3:8\to \cfrac{3}{8}\\ 12:32\to \cfrac{12}{32} \end{cases}\qquad \cfrac{\cancel{ 12 }}{\cancel{ 32 }}\implies \cfrac{\square ?}{\square ?} }\)

OpenStudy (anonymous):

 Cross multiply and the answer should be the same